# Revisiting Zeroth-Order Optimization for Memory-Efficient LLM Fine-Tuning: A Benchmark

### v2

@@ C How to Implement Memory-Efficient ZO/FO Optimizers?
$\underbrace{|{\mathbf{x}}|+|{\mathbf{s}}_{\text{opt}}|}_{\text{constant}}+\underbrace{\max\left\{|{\mathbf{s}}_{\text{fwd}}|,|{\mathbf{s}}_{\text{bwd}}|,|{\mathbf{s}}_{\text{opt}}^{\prime}|\right\}}_{\text{dynamic allocation}},$
where ||$\cdot$|| ⋅ | denotes the cardinality of a vector. Note that we concentrate on the primary factors in Big O notation, which, unless specified otherwise, was omitted for brevity.
C.1 Prevalent Efficiency-Enhancing Tricks Adopted in This Work
Prior to devling into the dissection of the memory consumption of different optimizers, we first introduce the prevalent efficiency-enhancing tricks widely adopted in the state-of-the-art FO/ZO-optimizers. These methods are by default adopted in this work if compatible to achieve the state-of-the-art performance.
Half-precision training (FP16). Most modern LLMs, e.g. LLaMA2 (Touvron et al., 2023), operate at half-precision (i.e., 16-bit) for inference, denoted as FP16. With FP16, only half of the model memory is required for ZO methods. By default, ZO methods are loaded at half-precision if the methods do not necessitate differentiation. However, as the current auto-differentiation module in PyTorch does not support FP16, methods like Forward-Grad must run at full precision. Similarly, for FO methods, half-precision is not compatible for the same reason.
Mixed-precision training (MP). Mixed precision is a common practice to speed up gradient computation by back-propagation. By default, MP is used in FO methods (FO-SGD and FO-Adam), where the model is loaded in full precision and the training is carried out in half-precision. Specifically, a half-precision replica of the model is used to conduct a forward pass yielding the intermediate results and gradients in half-precision. From the half-precision one, a full-precision gradient will be recovered for optimization. If the intermediate results consume more memory consumption than the model and gradients themselves, MP can also reduce memory complexity. We remark that MP is still operating on a full-precision model, which means MP and FP16 cannot be used together in PyTorch.
The ‘foreach’ implementation of Adam. The PyTorch implementation of Adam will use so-called foreach implementation to speed up the computation. At a high level, the foreach implementation will replicate and merge all the layers’ weight into one tensor during Adam updates. Operations on the unified tensor can be easily parallelized and therefore are faster. Though foreach can speed up computation, it demands extra memory to store all the weights.
C.2 Unpacking Memory Costs of Different Optimizers
In this section, we look into the optimizers studied in this work. For each algorithm, we will first analyze their theoretical memory consumption following the step $1\sim 3$ depicted in the framework in Algorithm A1. In the end, we provide a comparison between the theoretical and empirical results.
In most backward algorithms, activation ${\mathbf{a}}$ will be cached for backward computation. We use $\bar{\mathbf{a}}$ and $\bar{\mathbf{x}}$ to denote the 16-bit version of ${\mathbf{a}}$ and ${\mathbf{x}}$ (half-precision), respectively. For all FO methods, the memory of the model parameters is of size $|{\mathbf{x}}|$. For all ZO methods with FP16, the memory of parameters is halved as $\frac{1}{2}|{\mathbf{x}}|$. Next, we will focus on the memory dynamics in Step 1-3.
FO-SGD (MP). In Step 1, the forward pass will first replicate the model in half-precision, i.e. $\bar{\mathbf{x}}$, leading to an extra memory consumption $\frac{1}{2}|{\mathbf{x}}|$. Then, the activation per layer will be cached as half-precision intermediate results, $\bar{\mathbf{a}}=\oplus_{l=1}^{L}\bar{\mathbf{a}}_{l}$. In Step 2, the backward pass will compute the $16$-bit gradient using activation per layer. The computation happens layer by layer. Therefore, the peak memory will be summed as $\sum\nolimits_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}$. For simplicity, we approximate the term as $\max\{\frac{1}{2}|{\mathbf{a}}|,|{\mathbf{x}}|\}$. In Step 3, there will be no optimization state. We can approximate the dynamic memory as
$\quad\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\max\{\frac{1}{2}|{\mathbf{a}}|,|{\mathbf{x}}|\}\right\}$
$=\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,|{\mathbf{x}}|\right\}.$
The total memory consumption will be approximated as
$\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{3}{2}|{\mathbf{x}}|,2|{\mathbf{x}}|\right\}.$
FO-Adam (MP). In Step 1 and 2, FO-Adam completes the forward pass in the same way as FO-SGD and thus consumes memory similarly. In Step 3, the PyTorch Adam utilizes the ‘foreach’ implementation to speed up state computation. The Adam optimizer will first replicate the gradient in an extra memory of size $|{\mathbf{x}}|$. The replica will co-exist with the gradient in the memory temporarily, leading to a total memory $2|{\mathbf{x}}|$. Therefore, the peak memory will be the maximum of the dynamic memory in all of the three steps:
$\quad\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\max\{\frac{1}{2}|{\mathbf{a}}|,|{\mathbf{x}}|\},2|{\mathbf{x}}|\right\}$
$=\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,2|{\mathbf{x}}|\right\}.\vspace*{-3mm}$
In addition to the dynamic memory, the optimization step adds two states in memory for first-order and second-order momentum vectors, i.e., ${\mathbf{m}}_{t}$ and ${\mathbf{v}}_{t}$. We can sum up the memory of all steps to estimate the total memory as
$\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{7}{2}|{\mathbf{x}}|,5|{\mathbf{x}}|\right\}.\vspace*{-3mm}$
ZO-SGD (FP16). In Step 1-2, ZO-SGD estimates the gradient by two forward loss calculations based on the random direction vector ${\mathbf{u}}$. Yet, storing ${\mathbf{u}}$ will be as costful as the whole model parameter. To reduce the consumption, the per-layer perturbation ${\mathbf{u}}_{l}$ is computed on demand. For this purpose, we follow (Malladi et al., 2023) to use the random seed trick. Note the perturbation of all parameters is done sequentially and therefore we use random seed $\mathcal{S}$ to reproduce the random direction vectors in the same order. Specifically, we initialize a random number generator by $\texttt{rng}=\text{RandomState}(\mathcal{S})$. Sequentially, we compute the random vector for a specific parameter ${\mathbf{x}}_{l}$ by
${\mathbf{u}}_{l}\sim\mathcal{N}_{\text{rng}}(\mathbf{0},\mathbf{I}),$
where $\mathcal{N}_{\text{rng}}$ represent the Gaussian distribution with the generator rng. In Step 3, we can also reproduce the random vectors for computing the parameter gradient and immediately update ${\mathbf{x}}_{l}$. Without optimizer states, ZO-SGD consume a total memory as
$\frac{1}{2}|{\mathbf{x}}|+\max\nolimits_{l}|{\mathbf{u}}_{l}|=\frac{1}{2}|{\mathbf{x}}|+\max\nolimits_{l}|{\mathbf{x}}_{l}|,$
where the second term is the maximal dynamic memory with the random seed trick.
@@ C.2 Unpacking Memory Costs of Different Optimizers
ZO-SGD-MMT (FP16). ZO-SGD with Momentum is similar to ZO-SGD, but it consumes extra memory for momentum storage. The momentum shares the same size as the model parameter. According to the memory of ZO-SGD, we can get the memory estimation as
$|{\mathbf{x}}|+\max\nolimits_{l}|{\mathbf{x}}_{l}|.\vspace*{-5mm}$
ZO-Adam (FP16). Similar to ZO-SGD-MMT, ZO-Adam has extra optimizer states, first-order and second-order momentum. They have the same memory consumption as the parameters. Therefore, the total estimated memory is
$\frac{3}{2}|{\mathbf{x}}|+\max\nolimits_{l}|{\mathbf{x}}_{l}|.\vspace*{-5mm}$
Forward-Grad. In Step 1 and 2, Forward-Grad estimates the gradient in one unified pass. Different from other ZO methods, Forward-Grad leverages the forward-mode auto-differentiation module [^1] https://pytorch.org/tutorials/intermediate/forward_ad_usage.html in PyTorch to estimate the directional derivative, i.e., $\nabla f({\mathbf{x}})^{\top}{\mathbf{u}}$. Therefore, we have to create and maintain the random vector ${\mathbf{u}}$ for all parameters causing the memory in $|{\mathbf{x}}|$ size. In addition, Forward-Grad needs to forward the per-layer product of the Jacobian of activation w.r.t. the parameter and the random vector to the next layer. Therefore, there will be an additional memory cost per layer with the size of $|{\mathbf{a}}_{l}|$. In Step 3, Forward-Grad directly uses the gradient to update the parameters without extra memory costs. Thus, with the model memory, the total memory estimation can be summed as
$2|{\mathbf{x}}|+\max\nolimits_{l}|{\mathbf{a}}_{l}|.\vspace*{-4mm}$
Table A1: Comparison of total memory complexity of different optimizers when fine-tuning the full model. $|{\mathbf{x}}|$ denotes the memory of parameters (or gradients in the same size) in full precision. $|{\mathbf{a}}|$ denotes the memory consumption of intermediate results saved for post-hoc backward during forward. $|{\mathbf{x}}_{l}|$ and $|{\mathbf{a}}_{l}|$ represents the parameter and intermediate memory of a specific layer $l$.
Optimizer Memory
FO-SGD $|{\mathbf{x}}|+\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,|{\mathbf{x}}|\right\}$
FO-Adam $3|{\mathbf{x}}|+\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,2|{\mathbf{x}}|\right\}$
Forward-Grad $|{\mathbf{x}}|+|{\mathbf{x}}|+\max_{l}|{\mathbf{a}}_{l}|$
ZO-SGD $\frac{1}{2}|{\mathbf{x}}|+\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$
ZO-SGD-MMT $|{\mathbf{x}}|+\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$
ZO-Adam $\frac{3}{2}|{\mathbf{x}}|+\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$
C.3 Theoretical Memory Cost Comparison
In this section, we provide some analyisis based on the theoretical results discussed above. In particular, we first summarize the theoretical memory efficiency of different optimizers in Tab. A1. Several key insights can be summarized. First, compared to FO optimizers (including Forward-Grad), the memory efficiency merits of ZO optimizers are mainly in two aspects. On the one hand, ZO can avoid saving the intermediate results (model states) $\mathbf{a}_{l}$. As discussed in Fig. 2, this reduction is considerable when the sequence length of the language model is large. On the other hand, ZO method can estimate the gradient in a layer-wise sequential manner and thus avoid storing the gradients of the entire model. Secondly, in contrast to the half-precision (FP16) technique, the mixed precision (MP) training approach, which can be adopted by FO optimizers, solely diminishes the memory complexity associated with $|{\mathbf{a}}|$. It does not, however, reduce the memory consumption stemming from model loading or the storage of optimization states. In contrast, ZO can be equipped with FP16 with more flexibility, which helps achieves a memory reduction by $50\%$. Third, although Forward-Grad is also a BP-free method, its memory efficiency advantage over other FO optimizers is not remarkable. This is mainly due to its requirement for storing the intermediate results (i.e., $|{\mathbf{a}}_{l}|$). In the meantime, Forward-Grad is not compatible with the modern efficiency enhancing tricks (such as FP16 or MP), which further reduces its memory efficiency. Last, although ZO-Adam is memory intensive compared to ZO-SGD, it can be greatly improved by using FP16, which shows a better efficiency than FO-SGD (MP).
Theoretical analysis on the memory consumption change with sequence length. We remark that the empirical results ($|{\mathbf{a}}|$) in Tab. A1 are dependent on the input sequence length. In general, a larger sequence length will directly lead to a larger activation memory consumption. To investigate its impact, we explored how the scaling of input sequence length impacts memory usage by increasing activation memory in Fig. 2. First, when the sequence length is small (less than $700$), the total memory overhead remains approximately the same when the sequence length increase. This is primarily due to that the theoretical memory consumption for FO-SGD is $|{\mathbf{x}}|+\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,|{\mathbf{x}}|\right\}$ (see Tab. A1 for a reference) and the memory utilized by activations is less than that used for gradients. Second, as ZO-SGD does not need to store activations with a memory consumption of $\frac{1}{2}|{\mathbf{x}}|+\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$, it exhibits a much better efficiency advantage over FO-SGD, when the sequence length increases (e.g., $1504$). Therefore, the memory efficiency gap shown in Tab. 4 will be further enlarged if a larger input sequence length is used.

### v3

@@ C How to Implement Memory-Efficient ZO/FO Optimizers?
$\underbrace{|{\mathbf{x}}|+|{\mathbf{s}}_{\text{opt}}|}_{\text{constant}}+\underbrace{\max\left\{|{\mathbf{s}}_{\text{fwd}}|+|\bm{\tau}_{\text{fwd}}|,|{\mathbf{s}}_{\text{bwd}}|+|\bm{\tau}_{\text{bwd}}|,|{\mathbf{s}}_{\text{bwd}}|+|\bm{\tau}_{\text{opt}}|\right\}}_{\text{dynamic allocation}}.$
If the forward and backward are computed per layer and the temporary states can be voided, then the term $\max\{|{\mathbf{s}}_{\text{fwd}}|+|\bm{\tau}_{\text{fwd}}|,|{\mathbf{s}}_{\text{bwd}}|+|\bm{\tau}_{\text{bwd}}|\}$ will be replaced by $\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\}$, where $l$ represents any layer number, leading to a memory consumption as shown below:
$|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\},|{\mathbf{s}}_{\text{bwd}}|+|\bm{\tau}_{\text{opt}}|\right\}+|{\mathbf{s}}_{\text{opt}}|.$
In the next, we will go through all the FO/ZO optimizers considered in this work and discuss their memory efficiency one by one.
C.1 Theoretical Memory Efficiency Analysis of Different Optimizers
FO-SGD. Following Algorithm A1, the stored model/optimizer states are defined as
${\mathbf{s}}_{\text{fwd}}$ $=\oplus_{l=1}^{L}{\mathbf{a}}_{l}$
${\mathbf{s}}_{\text{bwd}}$ $=\oplus_{l=1}^{L}{\mathbf{g}}_{l}$
${\mathbf{s}}_{\text{opt}}$ ==$\emptyset$,= ∅ ,
where ${\mathbf{a}}_{l}$ represents the total activations being stored for computing the backward gradients ${\mathbf{g}}_{l}$, and $\oplus$ denotes the vector concatenation operation. Therefore, the total memory consumption is
$\quad|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\},|{\mathbf{s}}_{\text{bwd}}|\right\}+|{\mathbf{s}}_{\text{opt}}|$
$=|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{g}}_{l}|\},|{\mathbf{g}}|\right\}+0$
$=|{\mathbf{x}}|+\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{g}}_{l}|\}$
FO-SGD FP16. We use $\bar{\mathbf{a}}$ and $\bar{\mathbf{x}}$ to denote the 16-bit version of ${\mathbf{a}}$ and ${\mathbf{x}}$ (half-precision), respectively. The states are defined as:
${\mathbf{s}}_{\text{fwd}}$ $=\oplus_{l=1}^{L}\left(\bar{\mathbf{a}}_{l}\oplus\bar{\mathbf{x}}_{l}\right)$
${\mathbf{s}}_{\text{bwd}}$ $=\oplus_{l=1}^{L}{\mathbf{g}}_{l}$
${\mathbf{s}}_{\text{opt}}$ ==$\emptyset$.= ∅ .
Here, besides the activations $\bar{\mathbf{a}}_{l}$ stored in float16 for backward computing, there exists an extra copy of model weight in float16 in mixed precision training. This extra copy of model weight is a temporary memory, $|\tau_{\text{fwd}}|=|\bar{\mathbf{x}}|$, during forward computing. Therefore, the total memory consumption is
$\quad|{\mathbf{x}}|+\max\left\{|{\mathbf{s}}_{\text{fwd}}|+|\bm{\tau}_{\text{fwd}}|,|{\mathbf{s}}_{\text{bwd}}|+|\bm{\tau}_{\text{bwd}}|,|{\mathbf{s}}_{\text{bwd}}|\right\}+|{\mathbf{s}}_{\text{opt}}|$
$=|{\mathbf{x}}|+\max\left\{|\bar{\mathbf{a}}|+|\bar{\mathbf{x}}|,\sum_{l}\max\{|\bar{\mathbf{a}}_{l}|,|{\mathbf{g}}_{l}|\}\right\}.$
Vanilla ZO-SGD. For the vanilla implementation of ZO-SGD, the forward states include the random vector ${\mathbf{z}}$ and projected gradient $\delta$, i.e.,
${\mathbf{s}}_{\text{fwd}}={\mathbf{z}}\oplus\delta$
The projected gradient is computed by differentiating two forward passes, where the random vectors are given by:
${\mathbf{z}}=[{\mathbf{z}}_{l}\sim\mathcal{N}]_{l=1}^{L},$
which have the same dimension as the model parameters ${\mathbf{x}}$. Then, the backward state is the gradients estimated by ${\mathbf{z}}$ and no optimizer state is involved, i.e.,
${\mathbf{s}}_{\text{bwd}}$ $={\mathbf{g}}$
${\mathbf{s}}_{\text{opt}}$ ==$\emptyset$.= ∅ .
Therefore, the total memory consumption is:
$\quad|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\},|{\mathbf{s}}_{\text{bwd}}|\right\}+|{\mathbf{s}}_{\text{opt}}|$
$=|{\mathbf{x}}|+\max\{|{\mathbf{x}}|+|{\mathbf{g}}|\}+0$
$=|{\mathbf{x}}|+|{\mathbf{g}}|.$
ZO-SGD w/ memory reduction trick (ZO-SGD). In this work, the ZO-SGD method is by default implemented with the random state trick outlined in MeZO (Malladi et al., 2023) to reduce the memory consumption. The key idea of this trick is to save the random seed used to generate the random vectors instead of directly saving the random vectors themselves. By manually set the random seed, the same corresponding random vectors for each layer ${\mathbf{z}}_{l}$ can be generated on demand and used for perturbation without any additional storage incurred:
$\texttt{rng}=\text{randomState}(\mathcal{S}),\ {\mathbf{z}}=[{\mathbf{z}}_{l}\sim\mathcal{N}_{\text{rng}}]_{l=1}^{L}$
where rng is a pseudo-random variable generator. Therefore, all the states needed are:
${\mathbf{s}}_{\text{fwd}}$ $=\oplus_{l=1}^{L}{\mathbf{x}}_{l}$
${\mathbf{s}}_{\text{bwd}}$ $=\oplus_{l=1}^{L}{\mathbf{g}}_{l}$
${\mathbf{s}}_{\text{opt}}$ ==$\emptyset$,= ∅ ,
and this yields the total memory consumption:
$\quad|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\},|{\mathbf{s}}_{\text{bwd}}|\right\}+|{\mathbf{s}}_{\text{opt}}|$
$=|{\mathbf{x}}|+\max\{{\mathbf{g}}_{l}\}$
For simplicity, we use the name ZO-SGD to denote ZO-SGD w/ memory reduction trick throughout the paper.
@@ C.1 Theoretical Memory Efficiency Analysis of Different Optimizers
ZO-SGD-Momentum. ZO-SGD with Momentum is similar to ZO-SGD, but it consumes extra memory for momentum storage, which shares the same size as the model parameter. This is considered as an optimizer state in Algorithm A1. Therefore, the states are defined as:
${\mathbf{s}}_{\text{fwd}}$ $=\oplus_{l=1}^{L}{\mathbf{x}}_{l}$
${\mathbf{s}}_{\text{bwd}}$ $=\oplus_{l=1}^{L}{\mathbf{g}}_{l}$
${\mathbf{s}}_{\text{opt}}$ $=\oplus_{l=1}^{L}{\mathbf{x}}_{l}.$
Similar to ZO-SGD, the total memory consumption is:
$\quad|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\},|{\mathbf{s}}_{\text{bwd}}|\right\}+|{\mathbf{s}}_{\text{opt}}|$
$=|{\mathbf{x}}|+\max\{{\mathbf{g}}_{l}\}+|{\mathbf{x}}|$
$=2|{\mathbf{x}}|+\max\{{\mathbf{g}}_{l}\}.$
Forward Gradient. For FG, the forward states include the random vector ${\mathbf{z}}$ and projected gradient $\delta$. The projected gradient is computed by the Jacobian-Vector Product (JVP) and forward gradient, i.e., $\frac{\partial^{\top}f}{\partial{\mathbf{x}}}{\mathbf{z}}$, where the random vector is:
${\mathbf{z}}=[{\mathbf{z}}_{l}\sim\mathcal{N}]_{l=1}^{L}.$
In addition to ${\mathbf{z}}$ and $\delta$, JVP itself needs to store intermediate results. For example, consider a simplified two-layer network ${\mathbf{y}}=f_{2}(f_{1}({\mathbf{x}}))$. Suppose the output of the first layer is ${\mathbf{a}}=f_{1}({\mathbf{x}})$. The corresponding JVP can be formulated as
$\frac{\partial{\mathbf{y}}}{\partial{\mathbf{x}}}^{\top}{\mathbf{v}}$ $=\frac{\partial{\mathbf{y}}}{\partial{\mathbf{a}}}^{\top}\left(\frac{\partial{\mathbf{a}}}{\partial{\mathbf{x}}}^{\top}{\mathbf{v}}\right),$
where $\left(\frac{\partial{\mathbf{a}}}{\partial{\mathbf{x}}}^{\top}{\mathbf{v}}\right)$ is computed in the first layer and needs to take the memory of size $|{\mathbf{a}}|$. Now, we can summarize the memory required for the forward state ${\mathbf{s}}_{\text{fwd}}$ as
${\mathbf{s}}_{\text{fwd}}={\mathbf{z}}\oplus{\mathbf{a}}\oplus\delta.$
Then, the backward state is the gradients estimated by ${\mathbf{z}}$, i.e.,
${\mathbf{s}}_{\text{bwd}}$ $=\oplus_{l=1}^{L}\left(\delta{\mathbf{z}}_{l}\right)$
${\mathbf{s}}_{\text{opt}}$ ==$\emptyset$.= ∅ .
Therefore, the total memory consumption is
$\quad|{\mathbf{x}}|+\max\left\{\sum_{l}\max\{|{\mathbf{s}}_{\text{fwd},l}|,|{\mathbf{s}}_{\text{bwd},l}|\},|{\mathbf{s}}_{\text{bwd}}|\right\}+|{\mathbf{s}}_{\text{opt}}|$
$=|{\mathbf{x}}|+\max\left\{{\mathbf{g}}+\max\nolimits_{l}{\mathbf{a}}_{l},{\mathbf{g}}\right\}+0$
$=|{\mathbf{x}}|+|{\mathbf{g}}|+\max\nolimits_{l}\{{\mathbf{a}}_{l}\}.$
Forward Gradient (FG) without State is not supported in PyTorch. This is because PyTorch requires the ${\mathbf{z}}$ to be pre-computed, thus the ${\mathbf{s}}_{\text{fwd}}$ cannot be reduced to stateless implementation. Therefore, there is no implementation if we are using the forward gradient provided by PyTorch forward-mode automatic differentiation (forward-mode AD).
Though not possible with the PyTorch built-in Automatic Differentiation tools, FG without State is still possible. Specifically, for each layer, we keep the order when the ${\mathbf{z}}_{l}$ is computed, then the ${\mathbf{z}}_{l}$ can be computed on demand. The pursuit of a more memory-efficient FG computation in practice remains an open question, reserved for future work.
C.2 Other Implementation Details
Half precision (F16). Most modern LLMs are served at 16-bit float precision. By default, models will be loaded at full precision, i.e., 32-bit. To speed up inference and improve memory efficiency, models can be loaded at 16-bit precision, namely F16. We do F16 for ZO methods that do not require differentiation. Note for Forward Gradient, the F16 cannot be loaded for auto differentiation. For FO methods, the half-precision is not allowed for the same reason.
Mixed precision (FP16). For FO methods, mixed precision is a common practice to speed up gradient computation and reduce memory complexity. We denote the mixed precision (i.e., 16-bit on computing gradients) as FP16. Note that FP16 will not reduce the memory consumption for storing models or gradients but only affect the gradient computation and intermediate results.
For FO-SGD (FP16), the memory consumption for computing gradient is
$\max\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}\},$
where $\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|$ is the memory for model and activations in 16 bit and $\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}$ is the memory of activation and full-precision gradient.
The ‘foreach’ implementation of Adam. The PyTorch implementation of Adam will use the so-called ‘foreach’ implementation to speed up the computation. At a high level, the foreach implementation will merge all the layers’ weight into one tensor during Adam updates. Though foreach can speed up computation, it will demand extra memory to store all the weights.
Table A1: Memory consumption comparison for full fine-tuning (FT) evaluated on OPT-13B model with the MultiRC dataset (400 tokens per example on average). Notations are consistent with Tab. 5. The theoretical and empirical values below can be mutually corroborated, where $|{\mathbf{x}}|\approx 48$ GB, $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}\approx 49$ GB. Note $|{\mathbf{a}}|$ are dependent on the sequence length of the input.
Optimizer Theoretical Mem. Empirical Mem.
FO-SGD $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}+|{\mathbf{x}}|$ $97$ GB
FO-Adam w/o fast foreach $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}+3|{\mathbf{x}}|$ $195$ GB
FO-Adam $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}+4|{\mathbf{x}}|$ $239$ GB
Forward Grad $2|{\mathbf{x}}|+\max_{l}|{\mathbf{a}}_{l}|$ $103$ GB
Vanilla ZO-SGD $2|{\mathbf{x}}|$ $96$ GB
ZO-SGD $\max_{l}|{\mathbf{x}}_{l}|+|{\mathbf{x}}|$ $51$ GB
ZO-SGD MMT $\max_{l}|{\mathbf{x}}_{l}|+2|{\mathbf{x}}|$ $100$ GB
ZO-Adam $\max_{l}|{\mathbf{x}}_{l}|+3|{\mathbf{x}}|$ $151$ GB
FO-SGD (FP16) $\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}\right\}+|{\mathbf{x}}|$ $98$ GB
FO-Adam (FP16) $\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}\right\}+4|{\mathbf{x}}|$ $239$ GB
ZO-SGD (F16) $\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|+\frac{1}{2}|{\mathbf{x}}|$ $25$ GB
ZO-SGD-MMT (F16) $\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|+|{\mathbf{x}}|$ $49$ GB
ZO-Adam (F16) $\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|+\frac{3}{2}|{\mathbf{x}}|$ $74$ GB
Table A2: Memory consumption comparison for LoRA fine-tuning evaluated on OPT-13B model with the MultiRC dataset (400 tokens per example on average). Other settings are consistent with Tab. A1. The memory consumption of the LoRA adapter’s parameters is assumed to be significantly smaller than $|{\mathbf{x}}|$ and $|{\mathbf{a}}|$, and thus is omitted in this table.
Optimizer Theoretical Mem. Empirical Mem.
FO-SGD $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{\epsilon}}_{l}|\}+|{\mathbf{x}}|$ $69$ GB
FO-Adam w/o fast foreach $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{\epsilon}}_{l}|\}+2|{\mathbf{\epsilon}}|+|{\mathbf{x}}|$ $69$ GB
FO-Adam $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{\epsilon}}_{l}|\}+3|{\mathbf{\epsilon}}|+|{\mathbf{x}}|$ $69$ GB
Forward Grad $|{\mathbf{\epsilon}}|+|{\mathbf{x}}|+\max_{l}|{\mathbf{a}}_{l}|$ $55$ GB
Vanilla ZO-SGD $|{\mathbf{x}}|+|{\mathbf{\epsilon}}|$ $52$ GB
ZO-SGD $\max_{l}|{\mathbf{\epsilon}}_{l}|+|{\mathbf{x}}|$ $52$ GB
ZO-SGD MMT $\max_{l}|{\mathbf{\epsilon}}_{l}|+|{\mathbf{\epsilon}}|+|{\mathbf{x}}|$ $52$ GB
ZO-Adam $\max_{l}|{\mathbf{\epsilon}}_{l}|+2|{\mathbf{\epsilon}}|+|{\mathbf{x}}|$ $52$ GB
FO-SGD (FP16) $\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{\epsilon}}_{l}|\}\right\}+|{\mathbf{x}}|$ $92$ GB
FO-Adam (FP16) $\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{\epsilon}}_{l}|\}\right\}+3|{\mathbf{\epsilon}}|+|{\mathbf{x}}|$ $93$ GB
ZO-SGD (F16) $\max_{l}\frac{1}{2}|{\mathbf{\epsilon}}_{l}|+\frac{1}{2}|{\mathbf{x}}|$ $25$ GB
ZO-SGD-MMT (F16) $\max_{l}\frac{1}{2}|{\mathbf{\epsilon}}_{l}|+\frac{1}{2}|{\mathbf{\epsilon}}|+\frac{1}{2}|{\mathbf{x}}|$ $25$ GB
ZO-Adam (F16) $\max_{l}\frac{1}{2}|{\mathbf{\epsilon}}_{l}|+|{\mathbf{\epsilon}}|+\frac{1}{2}|{\mathbf{x}}|$ $25$ GB
C.3 Additional Experiments
We compare the empirical memory costs of full fine-tuning in Tab. A1 and LoRA fine-tuning in Tab. A2. Notably, the empirical results are generally aligned with the theoretical analysis. ZO methods are generally much more efficient than their FO counterparts, and the overhead can be further reduced by loading the half-precision model (using F16).